### v2

Abstract
In recent years, in an effort to promote fairness in the election process, a wide variety of techniques and metrics have been proposed to determine whether a map is a partisan gerrymander. The most accessible measures, requiring easily obtained data, are metrics such as the Mean-Median Difference, Efficiency Gap, Declination, and GEO metric. But for most of these metrics, researchers have struggled to describe, given no additional information, how a value of that metric on a single map indicates the presence or absence of gerrymandering.
Our main result is that each of these metrics is gameable when used as a single, isolated quantity to detect gerrymandering (or the lack thereof). That is, for each of the four metrics, we can find district plans for a given state with an extremely large number of Democratic-won (or Republican-won) districts while the metric value of that plan falls within a reasonable, predetermined bound. We do this by using a hill-climbing method to generate district plans that are constrained by the bounds on the metric but also maximize or nearly maximize the number of districts won by a party.
In addition, extreme values of the Mean-Median Difference do not necessarily correspond to maps with an extreme number of districts won. Thus, the Mean-Median Difference metric is particularly misleading, as it cannot distinguish more extreme maps from less extreme maps. The other metrics are more nuanced, but when assessed on an ensemble, none perform substantially differently from simply measuring number of districts won by a fixed party.
One clear consequence of these results is that they demonstrate the folly of specifying a priori bounds on a metric that a redistricting commission must meet in order to avoid gerrymandering.
keywords:
Metrics, Gerrymandering, Redistricting, Markov Chains, Hill Climbing
[…]
Acknowledgments This work is supported by the National Science Foundation under Grant No. DMS-1928930, while Ellen Veomett and Stephanie Somersille were in residence at the Mathematical Sciences Research Institute in Berkeley, California, during the semester/year of Fall/2023.
[…]
Data Availability All data used to conduct the empirical analyses were obtained from the Metric Geometry Gerrymandering Group’s GitHub organization at https://github.com/mggg-states.
[…]
Code Availability We have made all Python code used to run the empirical analyses publicly available at the following GitHub repository: https://github.com/stem-redistricting/ShortBursts/tree/main.
1 Introduction
Many legislative bodies elect their members through single member districts where jurisdictions are divided up into districts and voters elect one representative for their district. Perhaps the most well-known representatives elected in this way in the United States are the members of the House of Representatives, but districts are also used to elect many state senators, city council members, and others. These representatives are elected with the expectation (hope?) that they will represent their communities’ interests.
The United States has a long history of partisan gerrymandering where one party draws the maps to gain a durable, lasting advantage. Examples range from the original use of the term gerrymander in Massachusetts in 1812 to particularly egregious instances in the last twenty years which have led to multiple lawsuits that have reached the Supreme Court [GillWhitford, RuchoCommonCause, LamoneBenisek]. While the Supreme Court ruled in [RuchoCommonCause] that partisan gerrymandering is non-justiciable by the federal courts [^1] A matter is non-justiciable by federal courts if it is a matter upon which the federal court is not capable or not allowed to give a ruling., it remains justiciable in state courts where some states prohibit partisan gerrymandering. Furthermore, racial demographics and partisan preference are often intertwined and used as stand-ins for each other [RacePolitics]. Hence, the prevention of partisan gerrymandering can be a way to improve access and opportunity for historically disadvantaged racial groups.
A naturally appealing idea is to develop a metric to quantify gerrymandering: If a map has a metric value within some pre-determined bounds, then the map is reasonable, and if the metric value falls outside these bounds, then the map is gerrymandered. The existence of such a metric is very attractive since it would give citizens a means to evaluate any map produced by the party in control of the redistricting process. In addition, the metric could provide guidance for independent redistricting commissions in states that have such bodies, such as Arizona, California, Colorado, and Michigan.
The purpose of this paper is to examine four such metrics that have been proposed in the last 10 years: the Mean-Median Difference, Efficiency Gap, Declination, and GEO metric. The first three metrics are ubiquitous in the realm of the study of gerrymandering. For example, they can all be found in the GerryChain library [GerryDetails], a Python library developed to implement a Markov Chain Monte Carlo process constructing an ensemble of potential redistricting maps that could be used in comparison to a proposed redistricting map. They can all be found in Dave’s Redistricting App [DRA] and the PlanScore site [PlanScore], which are sites that allow the public to evaluate potential redistricting maps that they create or evaluate pre-existing maps. And these metrics can all be found in the Princeton Gerrymandering Project’s report cards [PGP], which are publicly viewable assessments of redistricting maps, done by the Princeton Gerrymandering Project team. For all of these resources ([GerryDetails, DRA, PlanScore, PGP]), the metrics are touted as indicating more extreme gerrymanders when the metric values are more extreme (farther from 0). The Geography and Election Outcome (GEO) metric is a more recently proposed metric that uses both map data as well as partisan data [geo_paper].
Our first main result is that each of these metrics is gameable when used as a single, isolated quantity to detect gerrymandering. That is, for each of the four metrics, we can find district plans for a given state with an extremely large number of Democratic-won (or Republican-won) districts while the metric value of that plan falls within a reasonable, predetermined bound. Indeed, we can find a single map with an extremely large number of Democratic-won districts on which all four of the metrics fall within the reasonable range.
Our second main result is that the Mean-Median Difference cannot distinguish between more extreme maps and less extreme maps at all. We describe the results more precisely in the following sections, but the main takeaway is that it is unrealistic to expect that a problem with as many dimensions as partisan gerrymandering could be reduced to a single quantity between 1-1 and $1$. These metrics may have a role to play as part of a more computationally intensive ensemble analysis, but we should not place any weight on the value of a single metric without further context. And the Mean-Median Difference should not be used at all, not even as part of a more computationally intensive ensemble analysis.
We perform three different studies which allow us to conclude these two main results, described in Sections 1.1, 1.2, and 1.3.
1.1 Gameability Study: Finding Extreme Maps that aren’t Out of Bounds
We define a map as “extreme” using the most straightforward of terms: it results an extreme number of districts won by a given party, while still maintaining the basic redistricting criteria of population balance, contiguity and some measure of compactness. Thus, in order to determine whether a metric is “gameable,” we must determine if extreme maps (maps with a high number of districts won by a particular party) are achievable while keeping the metric within acceptable bounds. In the case of the Efficiency Gap, the creators were careful not to prescribe the “acceptable bounds”. However, they used a bound of $\pm 8\%$ for the bound between “presumptive validity and invalidity” in their paper introducing the metric [PartisanGerrymanderingEfficiencyGap]. For consistency, we use the corresponding values here for all of our metrics: $\pm 16\%$ for metrics whose range is $\left[-1,1\right]$ (the GEO Metric Ratio [^2] A value obtained using the two GEO metric scores, which we define later. and Declination), and $\pm 8\%$ for metrics whose range is $\left[-0.5,0.5\right]$ (the Efficiency Gap and Mean-Median Difference). While we needed to fix a consistent range of metric values in order to complete our Gameability Study, in Section LABEL:sec:conclusions we explain why this choice does not impact our ultimate conclusions.
In order to create these extreme maps, we utilize a hill-climbing method dubbed the “Short Burst” method, developed by Cannon et al [cannonShortBursts]. The Short Burst method uses small runs of a Markov Chain Monte Carlo process to find extreme maps. Cannon et al showed that this Short Burst method finds maps with extreme numbers of majority-minority districts more successfully than other typically used methods, such as biased Markov Chains [cannonShortBursts]. While Cannon et al applied the Short Burst method to find districting plans with large numbers of majority-minority districts, we use it to find districting plans with largest number of majority-Democratic (or majority-Republican) districts. In section 3 we describe this method.
In Section LABEL:subsec:results_metrics_restricted, we discuss our results on the kinds of extreme maps that are allowed by each of the four metrics we consider. As we will see, each of these metrics is gameable for at least some state/map.
1.2 Extreme Metric Values versus Extreme Maps
While our Gameability Study focuses on extreme maps that have metric values within an acceptable range, one might also ask the even more basic question: do more extreme maps correspond to more extreme metric scores? This question is related to Stephanopoulos and McGee’s description of the efficiency principle. McGee introduced the efficiency principle in [MeasuringEfficiency], and based on Stephanopoulos and McGhee’s explanation of the efficiency principle in [MeasureMetricDebate], Veomett [2018arXiv180105301V] fleshed out the definition of the eficiency principle to have two parts:
The efficiency principle states that a measure of partisan gerrymandering must indicate both
EP1
a greater advantage for (against) a party when the seat share for that party increases (decreases) without any corresponding increase (decrease) in its vote share
EP2
a greater advantage against (for) a party when the vote share increases (decreases) without any corresponding increase (decrease) in the seat share.
In our exploration of extreme maps (and in every Markov Chain Monte Carlo process that creates an ensemble of maps), the same partisan data is used on every map in order to compare the maps to each other. Thus, the question of whether more extreme maps correspond to more extreme metric scores is similar to asking whether a metric satisfies a more flexible version of EP1; one in which the most extreme maps are correlated to extreme metric values.
In order to address this correlation question, we again use the Short Burst method. Without any restrictions (other than the required population balance, contiguity and loose compactness constraints) we create short bursts seeking maps with a large number of Democratic-won districts (and then repeat the process searching for maps with a large number of Republican-won districts). From the extreme maps that are generated, we separate them into maps with the maximum number of districts won $M$. We also consider maps with M1,M2,M3,M-1,M-2,M-3, and M4M-4 districts won. For each of those numbers, we look at the range of values for each metric on those maps to see how that range changes (or does not change) as the number of districts won changes.
In Section LABEL:subsec:results_no_restriction, we discuss the range of metric values on maps with the maximum or near-maximum number of Democratic-won (or Republican-won) districts. We consider the same metrics that we considered for the Gameability Study of Section 1.1: the Mean-Median Difference, Efficiency Gap, Declination, and GEO metric. We shall see that extreme values of the Mean-Median Difference do not correspond to maps with an extreme number of districts won (although extreme values of the other metrics do correspond to maps with an extreme number of districts won).
1.3 Metrics and Ensembles
Our third and final study is based in ensemble analysis. Ensemble analysis is a well-studied, widely agreed-upon technique in the modern study of gerrymandering. Groups from the Metric Geometry Gerrymandering Group [mgggWebpage] to the Quantifying Gerrymandering group [quantifyingGerrymanderingWebpage] to individual expert witness testimony [SS_PSC_MT] use ensemble analysis as a tool to detect gerrymandering.
Ensemble analysis is performed by creating a large number (an ensemble) of potential redistricting maps, which are expected to be a reasonable sample from the space of all redistricting maps. The proposed map is compared to the maps in the ensemble in some way, such as the number of districts won by a particular party, the order statistics of the vote shares in the districts, or the value of some other metric on those maps. If the proposed map looks like an outlier among the sample of maps, this is seen as indication that it was potentially drawn with partisan intent. [^3] This is the typical application of ensemble analysis, which has substantial support/usage in academic articles [DukeNC, DukeWisconsin, RecomMGGG, ColoradoInContext] and expert witness reports [EW_ClarkevWisconsin].
Using ensemble analysis on “number of districts won” by a particular party is a straightforward comparison when it comes to gerrymandering: we are determining whether the proposed map has an extreme number of districts won, compared to the other maps in the ensemble. The implications are easy to interpret. However, if we use ensemble analysis on another metric (such as the Mean-Median Difference, Efficiency Gap, Declination, and GEO metric) it is less clear how to interpret the results. Does an outlier Efficiency Gap value imply the map has an extreme number of districts won by some party? What about an outlier Mean-Median Difference value? Our results of this study are in Section LABEL:subsec:ensembles. In short, we see that the answer is “no” for the Mean-Median Difference (that is, outlier values of the Mean-Median Difference do not correspond to maps with an extreme number of districts won) and “yes” for the other metrics.
2 Metrics and Detecting Gerrymandering
As stated in Section 1, we will study four metrics which do not require detailed map data and are commonly used to detect partisan gerrymandering: the Mean-Median Difference, Efficiency Gap, Declination, and Geography and Election Outcome (GEO) metric. Each of these metrics assumes that there are two political parties vying for seats. We study these metrics both in terms of their gameability (results in Section LABEL:subsec:results_metrics_restricted) as well as whether extreme metric values correspond to extreme maps (results in Section LABEL:subsec:results_no_restriction). We also study how each of these metrics does or does not detect gerrymandering when evaluated on an ensemble of redistricting maps (results in Section LABEL:subsec:ensembles). For the reader unfamiliar with these metrics, we define them here. But first we highlight why we ground our understanding of the existance of gerrymandering in “districts won.”
2.1 “Districts Won” as the Foundation of Gerrymandering
In our Gameability Study, we seek to determine whether a metric is “gameable” by determining if a metric can be within reasonable bounds, but allow a map with an extreme number of districts won by some party. In our Extreme Metric Values versus Extreme Maps Study, we look to see if extreme values of each metric correlate with “extreme maps” as defined in Section 1.1: an extreme number of districts won by some party. And in our study of Metrics and Ensembles, we consider a map to be likely gerrymandered if that map’s number of districts won is an outlier among maps in the ensembles (that is, if the party has an extreme number of districts won).
Thus, for each of these studies, we are grounding our understanding of gerrymandering in an extreme number of districts won (or predicted to be won by partisan data). This certainly follows the common conception of gerrymandering. For example, the Brennan Center for Justice, in their “Gerrymandering Explained” report [BrennanCenterGerrymanderingExplained], describe gerrymanders in terms of the number of additional seats a party won due to gerrymandering. Intuition and common understanding aside, Political Scientists and Social Scientists also describe gerrymandering in terms of seats won: “Partisan gerrymanderers use their knowledge of voter preferences and their ability to draw favorable redistricting plans to maximize their party’s seat share [626161].” Mathematicians and Computer Scientists agree with this definition: “[T]he express intent of a partisan gerrymander is to secure for their party as many seats as possible under the constraints of voter geography and the other rules of redistricting” [ImplementingPartisanSymmetry]. While there may be other factors that appear in a gerrymandered map, such as durability of the map over time or incumbency protection, we hold that the baseline goal of a gerrymander is to push the number of districts won towards extreme values for the favored party.
2.2 Mean-Median Difference
The Mean-Median Difference is a partisan symmetry metric [ImplementingPartisanSymmetry] which was highlighted by MacDonald and Best in a 2015 article [doi:10.1089/elj.2015.0358]. This metric is calculated from partisan data only. Specifically, suppose a state has $n$ districts, whose estimated vote share for party $A$ is $V_{i}$ in district $i$ (such an estimation can be made using prior election data or any other partisan index estimating support for the two parties). The Mean-Median Difference is the median of the $V_{i}$s, minus the mean:
$\text{median}\left\{V_{i}:i=1,2,\dots,n\right\}-\text{mean}\left\{V_{i}:i=1,2,\dots,n\right\}$
A large positive Mean-Median Difference value indicates a larger median than mean, and the proponents of the Mean-Median Difference suggest that it indicates a map that is beneficial for party $A$.
The Mean-Median Difference is called a partisan symmetry metric because, under the assumption of “uniform partisan swing” (that when voters swing towards some party, they do so uniformly across all districts), the Mean-Median Difference measures an asymmetry in the way that the map treats the two parties. Specifically, it measures how far party $A$ can fall below a 50% vote share, while still receiving a 50% seat share [ImplementingPartisanSymmetry] [^4] To give a simple example, if party $A$ has a median vote share among the districts of $50\%$ and a mean vote share of $40\%$ (or a Mean-Median Difference of $0.1$) then party $A$ wins half the seats with only $40\%$ of the total vote share.. Thus, the Mean-Median Difference is always between -0.5 and 0.5.
2.3 Efficiency Gap
The Efficiency Gap is also a metric which is calculated from partisan data only. It was proposed by Stephanopoulos and McGhee [PartisanGerrymanderingEfficiencyGap] and was used to argue the existence of gerrymandering in Wisconsin in the Supreme Court case Gill v. Whitford [GillWhitford].
The Efficiency Gap utilizes the idea of a “wasted vote555“Wasted votes” is terminology from Stephanopoulos and McGhee’s paper, and not of our creation..” For a political party whose candidate lost in a district, all votes for that candidate are considered “wasted votes” for that party. For a political party whose candidate who won in a district, that party’s number of wasted votes from that district corresponds to the number of votes their candidate received beyond 50% of the votes in that district. That is, a winning candidate’s party’s wasted votes is the number of votes that candidate received, minus half of the turnout of that district.
To calculate the Efficiency Gap (EG), one sums party $B$’s wasted votes across all districts, subtracts the sum of party $A$’s wasted votes across all districts, and divides by the total votes:
[…]
A large positive EG value indicates more wasted votes for party $B$, which is intended to indicate a map that is beneficial for party $A$.
Note that exactly half of the votes in each district are wasted by one party or another. Thus, $0.5\leq EG\leq 0.5$.
2.4 Declination
The Declination is a metric introduced by Warrington, which also uses only partisan data [WarringtonDeclinationELJ]. This metric is intended to find “packing” and “cracking,” which are how maps are frequenty described to be gerrymandered. Packing and cracking occur when an opposing party’s voters are “packed” into a small number of districts that they win with an overwhelming majority. The remaining opposing party’s voters are “cracked” or split among many districts, so that they lose by relatively slim margins in those districts. It is by this packing and cracking that a mapmaker can attempt to win as many districts as possible for her preferred party.
To calculate the Declination, we order the vote shares $V_{i}$ in all districts $i=1,2,\dots,n$ from smallest vote share for party $B$ to largest vote share for party $B$:
$V_{1}\leq V_{2}\leq\cdots\leq V_{k}<\frac{1}{2}<V_{k+1}\leq\cdots\leq V_{n}$
Note that here we have party $B$ losing $k$ districts and winning nkn-k districts.
We then define two sets $\mathcal{A}$ and $\mathcal{B}$ as follows:
$\mathcal{A}$ $=\left\{\left(\frac{i}{n}-\frac{1}{2n},V_{i},\right):i=1,2,\dots,k\right\}$
$\mathcal{B}$ $=\left\{\left(\frac{j}{n}-\frac{1}{2n},V_{j}\right):j=k+1,k+2,\dots,n\right\}$
Let $F$ be the center of mass of points in $\mathcal{A}$, $G=\left(\frac{k}{n},\frac{1}{2}\right)$, and $H$ the center of mass of points in $\mathcal{B}$. We define angle $\theta_{P}$ to be the angle that the line through $F$ and $G$ makes with the line $y=\frac{1}{2}$, and $\theta_{Q}$ to be the angle that the line through $G$ and $H$ makes with the line $y=\frac{1}{2}$. These can be visualized in Figure 1, which is from from Warrington’s article [WarringtonDeclinationELJ]:
Figure 1: Image from [WarringtonDeclinationELJ], using data from North Carolina’s 2014 Congressional election
Then the declination $\delta$ is calculated as:
$\delta=\frac{2}{\pi}\left(\theta_{P}-\theta_{Q}\right)$
Note that the scaling by $\frac{2}{\pi}$ guarantees that $1\leq\delta\leq 1$. A large positive $\delta$ indicates that if the $V_{i}$’s correspond to party $B$’s vote shares, then party $B$’s “average” losing district666represented by point $F$ has a vote share near 50%, and party $B$’s “average” winning district777represented by point $H$ has a very high vote share (which translates into party $B$ being “packed and cracked”). Thus, a large positive $\delta$ is intended to indicate a map that is beneficial for party $A$.
2.5 Geography and Election Outcome (GEO) Metric
The GEO metric is the only metric we are aware of that can be computed by hand, and which also requires not just partisan data but also map data. It was proposed by Campisi, Ratliff, Somersille, and Veomett [geo_paper]. The GEO metric gives a score to each party which is a non-negative integer.
The calculation of party $X$’s GEO score involves swaps of vote shares between neighboring districts in a way that keeps any current wins for party $X$, but allows some lost districts to become competitive (have a 50/50 vote share split), without changing the vote shares so much as to be statistically unreasonable. The value of party $X$’s GEO score is a count: it is the number of additional districts that can be made competitive for party $X$ through this vote share swapping process (see [geo_paper] for more details). Thus, if party $X$’s GEO score is very large relative to the total number of districts, that party could potentially have many additional competitive districts with reasonable changes to the map (indicating the map is not favorable to party $X$). Conversely, if party $X$’s GEO score is close to 0, there is very little that can be done to improve party $X$’s chance of being competitive in additional districts (indicating the map is optimized for party $X$).
We note here that, for the Gameability Study and Extreme Metric Values versus Extreme Maps Study (described in Sections 1.1 and 1.2), we needed each metric to have a bounded range of values centered at 0 in order to compare “apples to apples” with respect to the values of the metrics. Thus, for those two studies, we used a “GEO ratio”, which is the difference of the two parties’ GEO scores, divided by the total number of districts in the state:
$\text{GEO ratio}=\frac{\text{GEO(B) - GEO(A)}}{\text{number of districts}}$
Since the GEO score for each party is an integer between 0 and the total number of districts, this GEO ratio is guaranteed to be between -1 and 1. The creators of the GEO metric do not recommend using the GEO ratio itself as a metric, and promote that the two GEO scores (along with the number of districts) be considered themselves when assessing a map [geo_paper]. Nevertheless, we included this GEO ratio so that we could include a variant of the GEO metric that gave a single score which was on a comparable scale to the other metrics. From the discussion above, we can see that a large positive GEO ratio indicates a map beneficial to party $A$.
3 Hill Climb Implementation: Short Bursts
There are many methods available to use in conjunction with a Markov Chain process to find plans that maximize a given metric. For example, one could accept a subsequent plan only when its metric value is greater than the last plan’s value. However there are many drawbacks to this method. It can be slow and computationally costly. Also, if the current position is a local maximum, one move in the chain to a higher valued plan may not be possible. Another commonly used method is a biased random walk. In a biased random walk, when moving to the next plan in the chain, a bias is imposed so that plans with higher values of the desired attribute are more likely to be chosen.
Cannon and her collaborators developed the Short Burst method as an alternative to a biased random walk [cannonShortBursts]. Specifically, it starts with a seed map and runs a Markov Chain Monte Carlo process for 10 steps (for example), to create 10 new redistricting maps. Among those maps, the map with the most extreme metric value is selected to be the seed of the next “short burst” of size 10; if two maps have the same extreme metric value, then the most recently visited map with that extreme metric value is retained as the next seed. This process is repeated thousands of times.888 Note that the Short Burst method can be thought of as a non-deterministic version of beam search on the metagraph of redistricting maps (edges are between two maps that can be obtained from each other through a single ReCom move in GerryChain), using depth-first search instead of breadth-first, and in which a random selection of maps near to the current map are explored.
In [cannonShortBursts], they then used the Short Burst method in a study of Louisiana’s state House of Representatives plans. Their goal was, in the context of a Voting Rights Act study, to attempt to find the maximum number of majority-minority districts a single districting plan could sustain. They found that, in the context of maximizing majority-minority districts, Short Bursts were more successful in finding extreme maps than biased random walks.
As stated in Section 1, we use the Short Burst method in order to create our “extreme” maps. i.e. the maps with the most Democratic-majority (Republican-majority) districts. Like Cannon et al, we use the freely available GerryChain library to implement the Markov Chain Monte Carlo (MCMC) process [GerryChain]. Based on the results in [cannonShortBursts], one can see that bursts of length 10 were able to find the most majority-minority districts, so we also used bursts of length 10 for our studies. In their study comparing short bursts to biased random walks (Section 5.2.3 of [cannonShortBursts]), the metric values drastically slowed in increase after about 5,000 steps. For this reason, and also because computing constraints, we limited our runs to 5,000 steps, completing 10 total runs (5,000 steps each) for each map. Given the fact that metric values continued to increase only slightly after 5,000 steps in the experiments from [cannonShortBursts], we expect that the number of districts won that we find in our experiments will be quite close to, but perhaps not equal to, the maximum possible number of districts won for that party and map. We used and heavily edited the code that Cannon’s team shared [SBGitHub]; our code is also available on GitHub at